OpenStudy (anonymous):

HHHHHEEEELLLPPPPPPPP!!!!!!!!!!!!!!!!!! A purse contains 21 coins consisting of nickels and dimes. Find the number of coins of each kind if the total value is $1.65. If 5n represents the value of the nickels, then which of the following expressions represents the value of the dimes? 10n 16n 10(21 - n)

OpenStudy (jdoe0001):

a nickel is 5cents a dime is 10cents, that is, twice as much of a nickel, or 2 nickels, or 2*nickels if a nickel = 5n then a dime is 2 * 5n ---------------------------------------------------- we know the purse has 21 coins, with only nickels and dimes say the quantity of nickels is "n" and the quantity of dimes is "d" so we could say that ---> n + d = 21 all nickels plus all dimes, add up to 21 coins total value of all is $1.65 we know that a nickel is 5cents and a dime is 10cents we also know that $1.65 is 165 cents (5 * quantity of nickels) + (10 + quantity of dimes) = 165 5n + 10d = 165 but a "d = 2n", twice as much as nickel 5n + (2n) = 165 # solve for "n" to find out how many nickel coins are there the remainder, will be dimes

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

5n + 10d = 165 but a "d = 2n", twice as much as nickel 5n + 10(2n) = 165

OpenStudy (jdoe0001):

lemme double-check that

OpenStudy (jdoe0001):

we know .... that all nickels and dimes add up to 21 coins and $165 pennies so \(\bf n+\color{red}{d}=21\\ \quad \\ 5n+10\color{red}{d} = 165\\ \quad \\ \quad \\ n+\color{red}{d}=21\implies d = \color{red}{21-n}\\ \quad \\ \textit{so we could say that}\\ 5n+10d = 165\implies 5n+10(\color{red}{21-n}) = 165\) solve for "n" to find out how many nickels there are the remaining change, is dimes :)

OpenStudy (anonymous):

thankx that helped...........